OpenStudy (anonymous):

gilda drove 650 miles at an average speed of 50 miles per hour. How many miles per hour faster would she have had to drive in order for the trip to have taken 1 hour less?

OpenStudy (ganpat):

speed = distance / time.. 50 = 650/ t so, t = 650 / 50 = 13 hrs.. to complete trip faster i.e. 1 hr less. i.e. in 12 hrs.. calculate speed.. sped = distance/ time.. speed = 650 / 12 = ?? can u find ??